# Look Twice Before You Answer: Memory-Space Visual Retracing for Hallucination Mitigation in Multimodal Large Language Models

### v1

@@ 4.2 Quantitative Results
Figure 9: Visualisation of uncertainty across layers without and with MemVR. MemVR effectively reduces uncertainty after 8th layer, contributing to hallucination mitigations as Assumption 3.1.
Q5: How does layer uncertainty assist with visual retracing?
MemVR leverages layer entropy to trigger visual retracing dynamically. We compared the performance of fixed retracing layers with dynamic retracing layers and found that dynamic retracing outperformed fixed retracing, as demonstrated in Table 4. Under all experimental conditions, dynamic retracing achieved the highest total score on the MME evaluation using LLaVA. Furthermore, we analyzed the injection ratio for MemVR, with the optimal injection ratio observed around 0.15, as shown in Table 4. This indicates that a balanced level of $\alpha$ is crucial for maximizing performance.
Q6: Can MemVR adapt to other MLLMs?
Definitely, MemVR is designed to be flexible and compatible with various architectures. A key advantage of MemVR lies in its ability to function without requiring modifications to the Transformers library, facilitating smooth integration into both older and cutting-edge models. We conducted extensive experiments across multiple benchmarks with LLaVA, Qwen-VL, and GLM-4V, and all three models outperformed baselines.
4.3 Qualitative Study: What types of faults can our method address?
@@
Figure 10: A case study in long text generation. MemVR effectively mitigates hallucinations.
In addition to evaluating single-word question-answering (QA) benchmarks, we further explore the models’ ability to generate comprehensive long-text descriptions for various tasks. As depicted in Figure 10, our MemVR excels in accurately identifying question-relevant details within images. In contrast, as discussed in Appendix C, Qwen-VL-Chat exhibits occasional difficulties in producing detailed image descriptions when applied to VCD. These shortcomings become evident in scenarios requiring a nuanced interpretation of image content. This suggests that MemVR demonstrates superior adaptability across different architectures, enabling more reliable long-text generation.
@@ 4.4 Limitations and Further Discussions
While MemVR demonstrates significant promise in improving the performance in MLLMs, it is not without limitations. One major challenge lies in the variability of MLLM architectures. Different MLLMs employ varying FNNs, activation functions, and knowledge systems, making it difficult to identify the optimal hyperparameters—such as injection ratios $\alpha$, and premature layer for each specific model. This requires considerable effort in model-specific tuning, which may limit the scalability of MemVR without further automation or standardization in hyperparameter selection.
Additionally, although our research focused on MLLMs with image inputs, MemVR is theoretically applicable to LLMs with different modalities. However, we have yet to explore its performance on inputs such as voice, point clouds, 3D meshes, or video frames. This opens up an exciting avenue for future work, where we plan to extend MemVR ’s framework to these diverse input formats and assess its efficacy across a broader range of tasks. Furthermore, we will explore integrating MemVR into the training procedures of MLLMs, rather than limiting its application to inference, to evaluate whether this could lead to even greater improvements in model performance and generalization.
@@ 5 Conclusion
This paper proposes a novel training-free paradigm to mitigate hallucination, named MemVR. In contrast to previous CD-based methods, which primarily focus on eliminating biases of language priors, MemVR seeks to replenish question-relevant visual clues towards more evidential answers, which signifies the other side of the coin. Our experiments, conducted on seven benchmarks, demonstrate the effectiveness of MemVR in mitigating hallucination and improving general performance.
    A: Ethic Considerations
  • Appendix
    B: Theoretical Analysis of MemVR in MLLMs
  • Appendix
    C: Additional Experiments, Results, and Discussions
Appendix A Ethic Considerations
We list some key ethical considerations of our method:
Bias and fairness. By injecting visual features from CLIP into the Feed-Forward Network (FFN) layers of LLMs, there’s a potential for inherited biases from the original models. CLIP, like many pre-trained models, may contain biases in how it represents certain objects, scenes, or demographics. These biases could propagate, affecting performance on different types of data (e.g., gender, ethnicity, cultural contexts). It’s important to evaluate how MemVR performs across diverse datasets and ensure that it doesn’t reinforce harmful stereotypes or disproportionately fail for certain groups.
Misuse of Enhanced Models. As MemVR aims to improve the long-text generation and overall performance of VLMs, enhanced models could be misused to generate deceptive or misleading content, such as deepfakes or disinformation. It’s important to consider whether there are safeguards in place to prevent malicious use of these improved models in scenarios like automated misinformation campaigns or unethical surveillance.
Data Privacy. If the benchmarks used for evaluating MemVR include datasets with personally identifiable information or sensitive content, care should be taken to ensure data privacy. Models should be evaluated on publicly available, anonymized, or ethically sourced datasets to avoid violating privacy laws or ethical norms.
Appendix B Theoretical Analysis of MemVR in MLLMs
In Multimodal Large Language Models (MLLMs), hallucinations often arise due to insufficient alignment between visual inputs and the model’s internal representations. This paper provides a rigorous theoretical analysis demonstrating that re-injecting visual features into the intermediate layers of MLLMs mitigates hallucinations and enhances representation capability.
@@ Appendix B Theoretical Analysis of MemVR in MLLMs
From Theorem B.2: $\hat{H}_{vq}^{(l)}$ contains more information about $Z_{v}$, i.e., $I(\hat{H}_{vq}^{(l)};Z_{v})\geq I(H_{vq}^{(l)};Z_{v})$. There is $I(Y;\hat{H}_{vq}^{(l)})\propto I(\hat{H}_{vq}^{(l)};Z_{v})$, thus we have $I(\hat{H}_{vq}^{(l)};Y)\geq I(H_{vq}^{(l)};Y)$. Then, we assume a dependency between $Z_{v}$ and $Y$, i.e., $I(Z_{v};Y)>0$, and subtract the inequalities, have:
@@
To ensure that $\mathcal{L}_{\Diamond}^{(m)}\leq\mathcal{L}^{(m)}$, we require: $\Delta\mathcal{L}\leq 0.$ We define the changes in mutual information. Let $\Delta I_{X}=I(H_{vq}^{(l)};X_{vq})-I(H_{vq}^{(l-1)};X_{vq})$, $\Delta I_{Y}=I(H_{vq}^{(l)};Y)-I(H_{vq}^{(l-1)};Y)$. Note that $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})\geq 0$. For $\Delta I_{X}$, the change in mutual information between $H_{vq}^{(l)}$ and $X_{vq}$ depends on how much additional information from $Z_{v}$ affects the dependence on $X_{vq}$. We denote the maximum possible increase as $\Delta I_{X}^{\max}$. For $\Delta I_{Y}$, From Theorem B.2, $\Delta I_{Y}\geq 0$, and suppose we can establish a minimum increase $\Delta I_{Y}^{\min}>0$. $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})$ represents supplement information about $Z_{v}$ in $H_{vq}^{(l)}$ that is not already explained by $X_{vq}$, and we denote this maximum as $I_{\max}^{Z\mid X}$.
@@
Appendix C Additional Experiments, Results, and Discussions
C.1 Benchmarks
@@
In this appendix, we provide additional details into the benchmarks referenced in the main paper. To evaluate hallucinations, we employ the following five benchmarks:
@@
CHAIR Rohrbach et al. (2018) evaluates how well the generated captions align with the content of the given image. CHAIR consists of two versions: CHAIR_S, which measures the inaccuracies at the sentence level, and CHAIR_I, which evaluates at the object level within the sentence by comparing the number of false objects to the total number of objects. For evaluation, we use the val2014 split of the MSCOCO Lin et al. (2014) dataset, which includes annotations for 80 object categories. We randomly select 500 images from the entire dataset and used the prompt “Please describe this image in detail.” for the MLLM.
@@
Polling based Object Probing Evaluation (POPE) Li et al. (2023c) is a VQA-based metric proposed to assess hallucinations in MLLMs. This metric evaluates the MLLM’s response to the prompt “Is [object] is in this image?” To emphasize that this is a binary VQA task, we appended the prompt with “Please answer yes or no.” To select objects referenced in the question prompt, we followed three different sampling options: random, popular, and adversarial. We evaluated performance across all sampling options.
@@
MLLM Evaluation (MME) Fu et al. (2023) evaluates the capabilities of MLLMs, dividing the evaluation into two major categories: perception and cognition. The perception category includes fine-grained tasks such as existence, count, location, rough color, poster, celebrity, scene, landmark, artwork identification, and OCR. The cognition category includes tasks like commonsense reasoning, numerical calculations, text translation, and code reasoning. All questions in this benchmark are structured to be answered with a simple yes or no.
@@
Using the LLaVA-Bench Liu et al. (2024), we further demonstrated how well our proposed method maintains the language model performance. This benchmark involves posing various situational questions, such as dialogue, detailed descriptions, and complex reasoning, to randomly selected images from the MSCOCO val2014 dataset. A total of 60 questions are used to assess whether the model faithfully follows the instructions. The generated answers are evaluated by comparing them to the responses of a text-only GPT-4 model.
C.2 Reproducibility
Implementation details.
@@ C.2 Reproducibility
We employed greedy search as the default decoding strategy across all benchmark evaluations. For the hallucination benchmarks (POPE and CHAIR) and general-purpose benchmarks (MME, VizWiz-VQA, MMBench, MM-Vet, and LLaVA-Bench (in-the-wild)), questions from the annotation files were used as prompts, formatted to fit the chat templates of each respective MLLM. Specifically, we utilized the COCO, A-OKVQA, and GQA datasets for POPE evaluation, and MMBench_DEV_EN for MMBench. In the MM-Vet evaluation, we used an online evaluator powered by OpenAI GPT-4 to assess generated results, while for LLaVA-Bench (in-the-wild), we employed OpenAI’s model gpt4-1106-preview via API. For CHAIR, a randomly sampled image set from the COCO Val2014 dataset was used across all three models, with the prompt ”Please describe this image in detail.” We sampled three different sets of images using different random seeds and evaluated performance by calculating the mean and standard deviation of the results. All MemVR tests were conducted using a greedy decoding approach, with do_sample=False, temperature=0, threshold=0.75, and beam=1. For VCD tests, we set do_sample=True, temperature=1, noise_step=500, and the plausibility constraint hyperparameter $\lambda$ to 0.1, while $\alpha$, which controls the degree of contrastive emphasis, was set to 1, following the default parameter settings from the original code and literature. OPERA tests were configured with beam=5, sample=True, scale_factor=50, threshold=15, num_attn_candidates=5, and penalty_weights=1. Due to OPERA’s reliance on older versions of Torch and Transformers, it was incompatible with Qwen and GLM models, and thus experiments involving these models were not conducted. Additionally, our method introduces two hyperparameters: the informative layer l for activation calculations and the factor $\lambda$ to control the influence of entropy on the next token probability distribution. To map the hidden states from selected layers l to vocabulary tokens, we chose intermediate layers based on the model’s depth (e.g., layers 5 to 16 for vicuna-7b, which has 32 layers), and we set $\lambda$ as a fixed value (e.g., 0.75). All parameter settings adhered to the default configurations specified in the respective papers and code repositories.
Experimental Code.
@@
To promote transparency and ensure the reproducibility of our work, we will release all experimental code, datasets, and detailed tutorials necessary for replicating our experiments. Our goal is to make it straightforward for researchers and practitioners to reproduce our results, regardless of their technical background. Additionally, by providing comprehensive documentation and clear guidelines, we aim to facilitate the extension of our method to other models and architectures, enabling the broader research community to explore its potential applications and improvements. We believe that open and reproducible research is essential for advancing the field and fostering collaboration.
Computational Resources.
@@
Our experiments were conducted on eight A40 and four A800 GPUs. The computational bottleneck was not the numerical accuracy values but the collection of potential hallucinatory factors for analytical purposes, including logits and attention values for each head and layer.
C.3 Case Study

### v2

@@ 5.4 Ablation Studies
Figure 8: Visualisation of uncertainty across layers without and with MemVR. MemVR effectively reduces uncertainty after 8th layer.
@@
Figure 9: A case study in long text generation. Compared with VCD, MemVR could effectively mitigate hallucinations.
Impact of VR Strategy. MemVR triggers VR dynamically using layer entropy. Tab. 10 shows that dynamic VR outperforms layer-fixed strategy, achieving the highest total score on MME benchmark, where Static-# denotes executing VR on the # layer and Static-$\phi$ means running VR on the specific layer with best performance. This indicates that dynamic VR leveraging layer entropy offers a more effective mechanism, adapting better to different scenarios and achieving optimal performance compared to static VR.
Table 10: MemVR performance with different VR strategies.
MME Static-7 Static-15 Static-23 Static-$\phi$ Dynamic
Cognition 347.1 352.14 357.9 362.9 383.9
Perception 1500.5 1529.1 1500.3 1526.4 1512.8
Total score 1847.6 1881.2 1858.1 1889.2 1896.7
5.5 Inference Latency
MemVR operates dynamically based on uncertainty. It uses VR when layer uncertainty exceeds threshold $\gamma$; if uncertainty stays low across all layers, indicating high model confidence, it’s not triggered. This mechanism enables efficient inference without extra computation. Different from CD-based and Att-intervention paradigms, which need multi-round inferences or have rollback-induced exponential overheads, our proposed look-twice mechanism only requires one regular inference. The comparisons on latency, throughput, time cost, and memory are shown in Table 1.
5.6 Cases Study
MemVR reduces uncertainty during inference. As shown in Figure 8, MemVR effectively reduces uncertainty after the 8th layer triggered VR, further confirming our analysis about the pattern of hallucinations in Section §3.3.
Long-text capability. Beyond single-word QA benchmark evaluation, we explore models’ capacity for comprehensive long-text generation in various tasks. As shown in Figure 9, MemVR can accurately identify image details relevant to questions. In contrast, as detailed in Appendix B, Qwen-VL-Chat struggles to generate detailed image descriptions in VCD, especially when nuanced image interpretation is needed. This indicates MemVR has better cross-architecture adaptability and more reliable long-text generation ability.
Analysis of failure cases. We have collected failure cases from the MME benchmark, in the ’Celebrity,’ ’Scene,’ and ’Landmark’ sub-tasks, where MemVR underperforms compared to the default model. We categorize MemVR’s failure cases into two types. Type 1 occurs when the default model is right but MemVR is wrong, due to over-disturbing the default model’s reasoning. In these instances, the original visual features are sufficient for reasoning, and the reinjected tokens inadvertently disrupt this process, leading to errors. We’re exploring ways to reduce such disturbances. For failure type 2, it happens when both are wrong, caused by image complexity or MLLM’s original knowledge gaps.
@@ 5.7 Limitations and Further Discussions
While MemVR shows substantial potential, it is not devoid of limitations. A primary challenge resides in the complex task of identifying the optimal hyperparameters. These hyperparameters include the injection ratio of visual information, and the strategy for selecting the triggered layers. This represents the focal point of our subsequent research efforts, as we work to perfect MemVR.
Additionally, although our research focused on MLLMs with visual inputs, theoretically, MemVR can be expanded to more modalities, e.g., listen-twice for audio, scan-twice for spatial perception, check-twice for fMRI, etc. This opens up an exciting avenue for future work, where we plan to extend MemVR’s framework to these diverse input formats and assess its efficacy across a broader range of tasks.
@@ 6 Conclusion
This paper proposes a novel decoding paradigm to mitigate hallucination, named MemVR, and present static and dynamic VR strategies that shift hidden states of the intermediate layer in MLLM for self-enhancement, rather than modulating logits directly with CD manner, thus avoiding multi-round decoding. Our experiments, conducted on eight benchmarks, demonstrate the effectiveness of MemVR in mitigating hallucination and improving general performance. Importantly, MemVR is a plug-and-play and task-agnostic method compatible with any MLLM, without extra fine-tuning, emphasizing its widespread applicability.
Acknowledgement
This work is sponsored by CAAI-Ant Group Research Fund; Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
Impact Statement
This paper proposes MemVR to alleviate hallucinations in MLLMs. The work has potential wider implications. Ethically, it may inherit biases from pre-trained models such as CLIP, risking unfair representation. There’s also concern that it could be misused to generate disinformation. In terms of societal impact, MemVR can improve system reliability in safety-critical areas such as healthcare.
Appendix A Theoretical Analysis of MemVR
In Multimodal Large Language Models, hallucinations often arise due to insufficient alignment between visual inputs and the model’s internal representations. This paper provides a rigorous theoretical analysis demonstrating that re-injecting visual features into the intermediate layers of MLLMs mitigates hallucinations and enhances representation capability.
@@ Appendix A Theoretical Analysis of MemVR
From Theorem A.2: $\hat{H}_{vq}^{(l)}$ contains more information about $Z_{v}$, i.e., $I(\hat{H}_{vq}^{(l)};Z_{v})\geq I(H_{vq}^{(l)};Z_{v})$. There is $I(Y;\hat{H}_{vq}^{(l)})\propto I(\hat{H}_{vq}^{(l)};Z_{v})$, thus we have $I(\hat{H}_{vq}^{(l)};Y)\geq I(H_{vq}^{(l)};Y)$. Then, we assume a dependency between $Z_{v}$ and $Y$, i.e., $I(Z_{v};Y)>0$, and subtract the inequalities, have:
@@
To ensure that $\mathcal{L}_{\Diamond}^{(m)}\leq\mathcal{L}^{(m)}$, we require: $\Delta\mathcal{L}\leq 0.$ We define the changes in mutual information. Let $\Delta I_{X}=I(H_{vq}^{(l)};X_{vq})-I(H_{vq}^{(l-1)};X_{vq})$, $\Delta I_{Y}=I(H_{vq}^{(l)};Y)-I(H_{vq}^{(l-1)};Y)$. Note that $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})\geq 0$. For $\Delta I_{X}$, the change in mutual information between $H_{vq}^{(l)}$ and $X_{vq}$ depends on how much additional information from $Z_{v}$ affects the dependence on $X_{vq}$. We denote the maximum possible increase as $\Delta I_{X}^{\max}$. For $\Delta I_{Y}$, From Theorem A.2, $\Delta I_{Y}\geq 0$, and suppose we can establish a minimum increase $\Delta I_{Y}^{\min}>0$. $I(H_{vq}^{(l)};Z_{v}\mid X_{vq})$ represents supplement information about $Z_{v}$ in $H_{vq}^{(l)}$ that is not already explained by $X_{vq}$, and we denote this maximum as $I_{\max}^{Z\mid X}$.
@@
Appendix B Additional Experiments, Results, and Discussions
B.1 Benchmarks and Metrics
Datasets. To rigorously assess the effectiveness of our proposed method, we conduct a comprehensive set of experiments across two benchmarks specifically designed to evaluate hallucination mitigation and five general-purpose benchmarks to gauge the general performance: ① Hallucination benchmarks: Polling-based Object Probing Evaluation (POPE) (Li et al., 2023b), and Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018), HallusionBench (Guan et al., 2024); ② General-purpose benchmarks: VizWiz-VQA (Gurari et al., 2018), MLLM Comprehensive Evaluation (MME) (Fu et al., 2023), Multimodal Benchmark (MMBench) (Liu et al., 2023b), Multimodal Veterinarian (MM-Vet) (Yu et al., 2024b), LLaVA-Bench (in-the-wild) (Liu et al., 2024b).
@@ B.1 Benchmarks and Metrics
In this appendix, we provide additional details on the benchmarks referenced in the main paper. To evaluate hallucinations, we employ the following five benchmarks:
@@
CHAIR (Rohrbach et al., 2018) evaluates how well the generated captions align with the content of the given image. CHAIR consists of two versions: CHAIRs, which measures the inaccuracies at the sentence level, and CHAIRi, which evaluates at the object level within the sentence by comparing the number of false objects to the total number of objects. For evaluation, we use the val2014 split of the MSCOCO (Lin et al., 2014) dataset, which includes annotations for 80 object categories. We randomly select 500 images from the entire dataset and used the prompt “Please describe this image in detail.” for the MLLM. The CHAIR metric includes per-instance evaluation (CHAIRi) and per-sentence evaluation (CHAIRs), defined as follows:
$\text{CHAIR}_{i}=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all objectsmentioned}\}|}$
$\text{CHAIR}_{s}=\frac{|\{\text{sentences with hallucinated object}\}|}{|\{\text{ all sentences}\}|}$
@@
Polling based Object Probing Evaluation (POPE) (Li et al., 2023b) is a VQA-based metric proposed to assess hallucinations in MLLMs. This metric evaluates the MLLM’s response to the prompt “Is [object] is in this image?” To emphasize that this is a binary VQA task, we appended the prompt with “Please answer yes or no.” To select objects referenced in the question prompt, we followed three different sampling options: random, popular, and adversarial. We evaluated performance across all sampling options.
@@
MLLM Evaluation (MME) (Fu et al., 2023) evaluates the capabilities of MLLMs, dividing the evaluation into two major categories: perception and cognition. The perception category includes fine-grained tasks such as existence, count, location, rough color, poster, celebrity, scene, landmark, artwork identification, and OCR. The cognition category includes tasks like commonsense reasoning, numerical calculations, text translation, and code reasoning. All questions in this benchmark are structured to be answered with a simple yes or no.
@@
Using the LLaVA-Bench (Liu et al., 2024b), we further demonstrated how well our proposed method maintains the language model performance. This benchmark involves posing various situational questions, such as dialogue, detailed descriptions, and complex reasoning, to randomly selected images from the MSCOCO val2014 dataset. A total of 60 questions are used to assess whether the model faithfully follows the instructions. The generated answers are evaluated by comparing them to the responses of a text-only GPT-4 model.
Candidate Layers. In dynamic premature layer selection, we partition transformer layers into buckets and select one bucket as the candidate layer set. For 32-layer LLaVA-1.5-7B, we use two buckets:[0,15),[15,31). This design limits the hyperparameter search space to only 2-4 validation runs. For efficiency, we use a validation set (MME) to select the best bucket.
B.2 Backbones and Baselines
To evaluate our method, we utilize well-known MLLMs: LLaVA-1.5 (Liu et al., 2024b), Qwen-VL (Bai et al., 2023), and GLM4V (Wang et al., 2023), LLaVA-Next (Li et al., 2024). Further, We compare our methods with classic training-tree SOTA methods designed to mitigate object hallucination, including visual contrastive decoding SOTA VCD (Leng et al., 2024), OPERA (Huang et al., 2024a) based on overconfidence penalty and hindsight allocation. As Dola (Chuang et al., 2023) is layer-wise contrastive decoding for LLMs and performs poorly in MLLMs, it will not be shown in the experiment. Experimental results are obtained and benchmarked using unified implementation.
Greedy search is used as the default decoding strategy in MemVR for all benchmarks. For benchmarks, annotation questions are adapted to MLLM templates. For POPE, COCO, A-OKVQA, and GQA are used, while MMBench_DEV_EN is used for MMBench. MM-Vet is assessed using MM-Vet Online Evaluator, and gpt4-1106-preview is used for LLaVA-Bench. CHAIR uses images from COCO Val2014 with the query ”Please describe this image in detail”. In MemVR, do_sample=False, temperature=0, threshold=0.75, beam=1. All settings of the compared method follow the default configurations from the original papers.
B.3 Reproducibility
@@
Implementation details. We employed greedy search as the default decoding strategy across all benchmark evaluations. For the hallucination benchmarks (POPE, CHAIR and HallusionBench) and general-purpose benchmarks (MME, VizWiz-VQA, MMBench, MM-Vet, and LLaVA-Bench (in-the-wild)), questions from the annotation files were used as prompts, formatted to fit the chat templates of each respective MLLM. Specifically, we utilized the COCO, A-OKVQA, and GQA datasets for POPE evaluation, and MMBench_DEV_EN for MMBench. In the MM-Vet evaluation, we used an online evaluator powered by OpenAI GPT-4 to assess generated results, while for LLaVA-Bench (in-the-wild) and HallusionBench, we employed OpenAI’s model gpt4-1106-preview and GPT-4o-mini respectively via API. For CHAIR, a randomly sampled image set from the COCO Val2014 dataset was used across all three models, with the prompt ”Please describe this image in detail.” We sampled three different sets of images using different random seeds and evaluated performance by calculating the mean and standard deviation of the results.
All MemVR tests were conducted using a greedy decoding approach, with do_sample=False, temperature=0, threshold=0.75, and beam=1. For VCD tests, we set do_sample=True, temperature=1, noise_step=500, and the plausibility constraint hyperparameter $\lambda$ to 0.1, while $\alpha$, which controls the degree of contrastive emphasis, was set to 1, following the default parameter settings from the original code and literature. OPERA tests were configured with beam=5, sample=True, scale_factor=50, threshold=15, num_attn_candidates=5, and penalty_weights=1. Due to OPERA’s reliance on older versions of Torch and Transformers, it was incompatible with Qwen and GLM models, and thus experiments involving these models were not conducted. Additionally, our method introduces two hyperparameters: the informative layer $l$ for activation calculations and the factor $\lambda$ to control the influence of entropy on the next token probability distribution. To map the hidden states from selected layers $l$ to vocabulary tokens, we chose intermediate layers based on the model’s depth (e.g., layers 5 to 16 for vicuna-7b, which has 32 layers), and we set $\lambda$ as a fixed value (e.g., 0.75). All parameter settings adhered to the default configurations specified in the respective papers and code repositories.
@@
Experimental Code. To promote transparency and ensure the reproducibility of our work, we will release all experimental code, datasets, and detailed tutorials necessary for replicating our experiments. Our goal is to make it straightforward for researchers and practitioners to reproduce our results, regardless of their technical background. Additionally, by providing comprehensive documentation and clear guidelines, we aim to facilitate the extension of our method to other models and architectures, enabling the broader research community to explore its potential applications and improvements. We believe that open and reproducible research is essential for advancing the field and fostering collaboration.
@@
Computational Resources. Our experiments were conducted on eight A40 and four A800 GPUs. The computational bottleneck was not the numerical accuracy values but the collection of potential hallucinatory factors for analytical purposes, including logits and attention values for each head and layer.
B.4 Case Study